OpenStudy (anonymous):

help to Factor: x2 - 100

OpenStudy (anonymous):

use a^2 - b^2 = (a-b)(a+b)

OpenStudy (anonymous):

i got it thanks :)

OpenStudy (amistre64):

(x+a)(x+b) = x^2 +bx + ax + ab = x^2 + (b+a)x + ab a and b are products of the last term that add up to the middle term the middle term is zero in this case